One number is four less than three times another. If their sum is increased by five, the result is twenty five. Find the numbers.
Answer by mananth(16949)

One number is four less than three times another. If their sum is increased by five, the result is twenty five. Find the numbers.
...
let the number be x
other number = 3x-4
..
x+3x-4+5 =25
4x+1=25
4x=25-1
4x=24
x= 24/4
x= 4
the other number = 3x-4
=12-4
=8
..
numbers are 4 & 8